OpenStudy (anonymous):

Find the slope of the line that passes through the points (6/7, 1 1/5) and (4/7, 2/5)

OpenStudy (anonymous):

A. 5/14 B. 8/35 C. 2 4/5 D. 4 7/35

OpenStudy (anonymous):

where are you stuck at? fractions? or the slope equation?

OpenStudy (anonymous):

I literally have no idea what to do @DemolisionWolf

OpenStudy (shamil98):

\[\frac{ y_2-y_1 }{ x_2-x_1 } = \frac{ \frac{ 2 }{ 5 } -\frac{ 6 }{ 5 } }{ \frac{ 4 }{ 7 } -\frac{ 6 }{ 7 } }\]

OpenStudy (anonymous):

I got 2.8

OpenStudy (anonymous):

so the slope of a line means 'rise over run' so how much you have to go up from one point to get to another is the 'rise' part, and the 'run' is hoe much you have to go sideways to get from one point to another. shamil98 has the slope equation written out.

OpenStudy (shamil98):

(2-6)/5 / (4-6)/(7) = (-4/5)/(-2/7) = \[\frac{ -4 }{ 5} \times \frac{ -7 }{ 2 }\]

OpenStudy (anonymous):

so it would be C?

OpenStudy (shamil98):

-4 x - 7 = 28 5 x 2 = 10 28 / 10 = 14/5

OpenStudy (shamil98):

14/5 = 2 4/5 so yes, C

OpenStudy (anonymous):

Thank you!